1 pt) A bacteria culture initially contains 2000 bacteria and doubles every half hour.
Find the size of the bacterial population after 100 minutes.
Find the size of the bacterial population after 6 hours.

Answer by stanbon(75887)

A bacteria culture initially contains 2000 bacteria and doubles every half hour.
---
A(t) = Ao(2^(t/30)) where t is minutes
--------------------------------------------
Find the size of the bacterial population after 100 minutes.
A(100) = 2000*2^(100/30) = 2000*10.079 = 20158.74 bacteria
-------------------------------
Find the size of the bacterial population after 6 hours.
6 hrs is 360 minutes
A(360) = 2000*2^(360/30) = 2000*2^12 = 8192000 bacteria
======================
Cheers,
Stan H.